OpenStudy (anonymous):

help pls:/ Sin(theta/2)=1/2 solve the following? 1. : Write an equation that expresses the value of theta/2 in terms of an appropriate inverse trigonometric expression 2. On the interval [0,2pi] what values of theta/2 satisfy your equation in number 1? 3. On the interval [0,2pi], what values of theta does the answer to number 2 produce? 4. Write an expression for all solutions to the equation

OpenStudy (mertsj):

\[\sin^{-1} (\frac{1}{2})=\frac{\theta}{2}\]

OpenStudy (mertsj):

\[\frac{\theta}{2}=\frac{\pi}{6}, \frac{5\pi}{6}\]

OpenStudy (mertsj):

\[\theta=\frac{\pi}{3}, \frac{5\pi}{3}\]

OpenStudy (mertsj):

\[\theta=\frac{\pi}{3}, \frac{5\pi}{3}+2\pi n\]

OpenStudy (anonymous):

(cosx-sqrt2/2)(secx-1)=0 use reciprocal identity to express the equation involving secant in term of sine,cosine,or tangent.